OpenStudy (mouth4war):

what is the volume of a right cone with a radius of 12 cm, height of 8 cm? (the side length is 14.4 cm if that makes a difference)

OpenStudy (campbell_st):

well the volume of a cone is \[V = \frac{1}{3} \times \pi \times r^2 \times h\] you have been given the radius and height, so just substitute and calculate the answer

OpenStudy (mouth4war):

i have done that, multiple times, but ive never gotten an answer thats part of the choices they gave me